9a.An artillery shell is fired with an initial velocity of 300 m/s at 62.5° above the horizontal. To clear an avalanche, it explodes on a mountainside 34.5 s after firing. What are the x- and y-coordinates of the shell where it explodes, relative to its firing point?


1
Expert's answer
2022-06-06T14:58:20-0400
"x=v_0\cos\theta t\\
=300*\cos62.5^\circ*34.5=4780\:\rm m"

"y=v_0\sin\theta t-gt^2/2\\
=300*\sin62.5^\circ*34.5-9.8*34.5^2/2=3350\:\rm m"